ABC Corp has a beta of 1.5, the current risk-free interest rate is 3%, and the expected return on the market is 12.0%. The market cap of ABC is $2 billion and the company has one 10-year bond outstanding with face value of $1 billion and coupon of 5%. The bond is trading at a premium price of 110. The company’s marginal tax rate is 25%. What is the ABC’s WACC?
Answer #1
D/A = D/(E+D)
D/A = 0.5/(1+0.5)
=0.3333
Weight of equity = 1-D/A
Weight of equity = 1-0.3333
W(E)=0.6667
Weight of debt = D/A
Weight of debt = 0.3333
W(D)=0.3333
Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (expected return on the market - risk-free rate)
Cost of equity% = 3 + 1.5 * (12 - 3)
Cost of equity% = 16.5
Cost of debt
                  K = N
Bond Price =∑ [(Annual Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =10
1100 =∑ [(5*1000/100)/(1 + YTM/100)^k]     +   1000/(1 + YTM/100)^10
                   k=1
YTM = 3.7805241351
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 3.7805241351*(1-0.25)
= 2.835393101325
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=2.84*0.3333+16.5*0.6667
WACC =11.95%